q1=1000-150p+2y+20p2 if current prices good 1 is 5 birr the price of good 2 is 8 birr and par capital of income is 500 birr where Q1is quantity of good 1 demand, p1 is prices of good 1, p2 is price of good 2,and Y is per capita income of the consumer then calculate prices elasticity of demand
1
Expert's answer
2021-08-16T08:56:36-0400

"q_1=1000-150p+2y+20p_2\\q_1=1000-150(5)+2(500)+20(160)\\=1000-750+1000+160\\=1410"

"e_{(p)}\frac{\frac{Q_1-Q_0}{(Q_1+Q_0)/2}}{\frac{P_1-P_0}{(P_1+P_0)/2}}"

ep=price elasticity

Q1=quantity of the demanded good 1

p1=price of the demanded good 1

Q0=initial quantity

p0=initial price

"e_{(p)}\frac{\frac{1410-0}{(1410+0)/2}}{\frac{5-0}{(5+0)/2}}\\=\frac{\frac{1410}{750}}{\frac{5}{2.5}}\\=\frac{2}{2}\\=1"